a car tavelling at 60km/h overtakes another car travelling at 42km/h.assuming each car to be 5m long, find the timetaken during the overtake and the total road distance?

Grade:Upto college level

1 Answers

gaurav bansal
11 years ago

first car velocity=42km/h second car=60km/h their relative velocity=18km/h=5m/s the car has to travel total distance=10m time taken=10/5s => 2s total distance covered=60*1000*2/3600 =50m